OpenStudy (tansheet):

a computer user tries to recall his password. she knows it can be one of 4 possible passwords. she tries her passwords until she finds the right one. let x be the number of wrong passwords she uses before she finds the right one. find E(x) and Var(x).

OpenStudy (ash2326):

She knows it's 1 of the 4 passwords x= no. of wrong password before she finds the right one she can guess correct at first so x=0 Probability of this \[P(x=0)= \frac{1}{4}\] she guesses correct at second time \[P(x=1)= P(getting\ wrong\ at\ first) \times P( getting\ right\ at\ second\ time)\] \[P(x=1)=\frac{3}{4} \times \frac{1}{3}\]

OpenStudy (ash2326):

Did you understand till this?? then we'll proceed

OpenStudy (tansheet):

what is math processing error...??

OpenStudy (ash2326):

Oh just reload the page

OpenStudy (tansheet):

understand till this.... you can procced ...?

OpenStudy (ash2326):

Just wait for sometime. I'll solve it

OpenStudy (ash2326):

So we have \[P(x=1)= \frac{1}{4}\] Now Let's find the probability of getting two wrong before getting right \[P(x=2)= \frac{3}{4} \frac{2}{3} \frac{1}{2}=\frac{1}{4}\] Probability of getting three wrongs \[P(x=3)= \frac{3}{4} \frac{2}{3} \frac{1}{2} \frac{1}{1}=\frac{1}{4}\] So we have x P(x) 0 1/4 1 1/4 2 1/4 3 1/4 \[E(x)=\sum_{x=0}^{x=3} x P(x)= 0 * 1/4+1*1/4+2*1/4+3*1/4=1.5\]

OpenStudy (ash2326):

Did you understand this?

OpenStudy (tansheet):

yes..... i do understand

OpenStudy (ash2326):

Great, now can you find variance?

OpenStudy (tansheet):

yes..

OpenStudy (ash2326):

Good :)